91 2.0 1.0 0 0.2 0.4 Volume, m3 0.6 0.8

A gas expands along path ABC. The work done by the gas in this expansion is 7.0 * 10^5 J

It shows part of a cycle using one mole of a monatomic, ideal gas as the working fluid.
Answer #1

92) Heat capacity at constant volume is given by:

Q = nc_{v}\bigtriangleup T

Given n=1 (no. of moles), and c_{v} = \frac{3}{2}R for a monoatomic ideal gas and heat capacity is defined for temperature change of \bigtriangleup T =1

So,

Heat\ capacity = \frac{3}{2}R

93) Heat capacity at constant pressure is given by:

Q = nc_{p}\bigtriangleup T

Given n=1 (no. of moles), and c_{p} = \frac{5}{2}R for a monoatomic ideal gas and heat capacity is defined for temperature change of \bigtriangleup T =1

So,

Heat\ capacity\ (constant\ pressure) = \frac{5}{2}R

94) change in internal energy is given by:

\bigtriangleup E_{int} = nc_{v}\bigtriangleup T

So from A to A, the pressure and volume of the gas will be same and so the tempereture acoording to ideal gas law, PV = nRT

So, \bigtriangleup T = 0

and hence \bigtriangleup E_{int} = 0

95) acoording to ideal gas law, PV = nRT

T \propto PV

So, \frac{T_{A}}{T_{C}} = \frac{P_{A}V_{A}}{P_{C}V_{C}}

\frac{T_{A}}{T_{C}} = \frac{2\times 10^{6}\times 0.4}{1\times 10^{6}\times 0.8}

\Rightarrow \frac{T_{A}}{T_{C}} = 1

\Rightarrow T_{A}= T_{C}

So, temperature at A and C will be same

96) Q at pressure = 2 * 106 ​Pascal will be

Q_{2} = nC_{p}\bigtriangleup T

\Rightarrow Q_{2} = nC_{p}\frac{\bigtriangleup (PV)}{nR}

\Rightarrow Q_{2} = n\times \frac{3}{2}R\times \frac{0.2\times 10^{6}}{nR}

\Rightarrow Q_{2} = 0.3\times10^{6}

Since, Q_{2} > 0​ So direction is inside the system